Please help me with this question!
y=-x+4
x+2y=12

Answer by mananth(16946)

x+y=4 .............1
x+2y=12 .............2
Eliminate y
multiply (1)by -2
Multiply (2) by 1
-2x-2y= -8
x+2y=12
Add the two equations
-x= 4
/-1
x=-4
plug value of x in (1)
x+y=4
-4+y=4
y=4+4
y=8
y=8